OpenStudy (lgbasallote):

how to integrate \[\int \frac{8}{4v^2 +1}dv\] lol i forgot how to do these

OpenStudy (anonymous):

Can try taking the constant, 8, out, and then do a backwards-chain rule operation that will leave you with a natural log expression..

OpenStudy (zarkon):

\[\int \frac{8}{4v^2 +1}dv\] \[=\int \frac{8}{(2v)^2 +1}dv\] Let\[2v=\tan(\theta)\]...

OpenStudy (lgbasallote):

*facepalm* tangent of course

OpenStudy (zarkon):

:)

OpenStudy (anonymous):

I had a feeling a trig-sub would work. Whenever you see a sum-of-squares, think trig-sub!

OpenStudy (lgbasallote):

is it 1/2 arctan 2v?

OpenStudy (zarkon):

close

OpenStudy (zarkon):

you forgot your 8 i believe

OpenStudy (zarkon):

\[4\,\tan^{-1}(2v)+c\]

OpenStudy (lgbasallote):

8?

OpenStudy (lgbasallote):

oh i see

OpenStudy (zarkon):

\[\int \frac{8}{4v^2 +1}dv=8\int \frac{1}{4v^2 +1}dv\] \[=8\frac{1}{2}\arctan(2v)+c=4\arctan(2v)+c\]